1.) Both total revenues (TR) and total costs (TC) are likely to be affected by changes...

1.)
Both total revenues (TR) and total costs (TC) are likely to be affected by changes in the output.
True or False
2.)
Profit is the unit contribution margin multiplied by the number of units minus the fixed component of the total costs (TC).
True or False
3.)
The average selling price is $.60 per unit, the average variable cost is $.36 per unit, and the total fixed costs are $1,500. If operating profits of $900 are desired, what is the sales volume in units?
4.)
The difference between total sales in dollars and total variable costs is called:
A.
operating profit.
B.
net profit.
C.
the gross margin.
D.
the contribution margin.

5.)
The following information pertains to Tiller Co.:
Sales
$800,000
Variable Costs
160,000
Fixed Costs
40,000

What is Tiller's break-even point in sales dollars?
Answer #1

Q1) It is True Statement

Q2)  It is True Statement

Q3) Let Sales volume be x

x(0.60-0.36)-1500 = 900

x = 10000

Q4) Option D. the contribution margin.

Q5) CM = (800000-160000)/800000 = 0.80

break-even point in sales dollars = 40000/0.80 = $50000

So, break-even point in sales dollars = $50000